Open Source Writers Group: Roadmap

1. Short-Term Development Plans

Currently, the OSWG has four mailing lists, a website, and a CVS tree. These are a good start towards establishing a foundation for the project, but there is definitely more work to be done.

1.1. Automate Documentation Generation

First off, the automatic generation scripts that turn the DocBook SGML instances contained in the CVS tree have to be completed. These scripts are currently (as of Nov 1, 1999) in the "testing" phase, and should be fully automated within a week or so. More scripts are needed, however, including db2man (for man pages) and db2ascii (for plaintext ascii). Current scripts (which are part of the docbook-tools suite) enable the autogeneration of HTML, PS, PDF, DVI, and TEX formats.

1.2. Improve Open-Content DocBook Docs

Second, the OSWG must increase the quantity and quality of it’s DocBook-related resources. Ideally, the OSWG will have DocBook documentation written by one or more OSWG members that includes templates, tutorials, various HOWTOs and FAQs, etc. The purpose of these resources would be to make it as easy as possible for new people to begin writing and submitting documentation in valid DocBook format.

Open Source Writers Group: Roadmap

1.3. Create a Proper Volunteers/Projects Database

Also on the OSWG TODO list is the creation and population of a volunteers/projects database that will automate the process of registering volunteers and projects and adding those to the OSWG website. The database will also provide some functionality towards automating the process of matching volunteers to projects and vice-versa. The "documentation wishlist" will also be included, as will all OSWG-initiated projects.

1.4. Improve and Add to Existing Communications Channels

In order to further open the channels of communication, it is hoped that the OSWG will be able to set up an IRC server (ideally irc.oswg.org) to host IRC channels for any documentation project that wishes to take advantage of this service. Naturally, #oswg will be the OSWG’s own channel. This shouldn’t be difficult to arrange.

1.5. Develop a Cooperative Relationship with the OSRT and Similar Groups

The OSWG could do valuable work with the Open Source Research Team that is part of MetaLab. The OSRT is doing some extremely interesting work regarding metadata and open-source documentation, and the OSWG could and should work with them to improve their ideas and implementations, as well as acting to educate and recruit people to help with the future iterations of this project.

2. Long-Term Development Plans

2.1. Open-Content Documentation Symposium/Conference

Farther down the road, the OSWG intends to organize and host an Open-Source/Open-Content Documentation Symposium. Initial planning for this event has already begun, and details will be announced as possible. The success and feasibility of the Symposium depend a great deal on the generosity of interested sponsors.

2.2. Open-Content Documentation Grant Program

Ideally, the OSWG will also be able to act as a central organization through which a documentation-grant program can be initiated. Money for this program would be donated by interested sponsors, and would be given to writers or project teams who have valuable projects (with solid project plans) that will further increase the quality and quantity of free open-source/open-content documentation. Details for the grant program must be discussed at great length, of course, as this idea is currently in the "gestation" phase.

2.3. Open-Content Documentation Awards

The OSWG also plans on initiating the Annual Open-Source/Open-Content Documentation Awards. These awards would be given to writers or project teams who have done outstanding work that contributes to the overall quality and quantity of o-s/o-c documentation. Again, this idea is currently in the development stages, and must be refined. The number and value of awards would be dictated by the generosity of interested sponsors.

3. Further Future Development

Future development of the OSWG and its projects will depend a great deal on its members as well as on the Open Source community as a whole. If you have any ideas or feedback about how the OSWG could better fulfill its primary goal of improving the quality and quantity of free open-souce/open-content documentation, please email editor@oswg.org. The OSWG is a community project – wholly for the community, and wholly by the community. It will endeavour to meet the documentation and related needs of that community, above all else.